OpenStudy (johndp123456):

pls help will fan and medal solve the following equation 6=x+2/3

OpenStudy (johndp123456):

@Awolflover1

Awolflover1 (awolflover1):

im here to help :)

Awolflover1 (awolflover1):

\[6=x+\frac{ 2 }{ 3 }\]

Awolflover1 (awolflover1):

lets flip it :) \[x+ \frac{ 2 }{ 3 } =6\]

Awolflover1 (awolflover1):

now Subtract 2/3 on both sides \[x+\frac{ 2 }{ 3 }-\frac{ 2 }{3}=6-\frac{ 2 }{ 3 }\]

OpenStudy (johndp123456):

16/3?

Awolflover1 (awolflover1):

Yess! There we go!

OpenStudy (johndp123456):

thanks

Awolflover1 (awolflover1):

you're welcome